Suppose the market for corn in Ghana is described by the following demand and supply

equations:

Demand: P = 100 – (1/2)Q

Supply: P = 10 + (13/10)Q.

 Where, P is price(₵/bag); and, Q is quantity(bags). Use this information to answer the

following questions.

a. What is the equilibrium price and quantity of corn in this market?

b. Plot the demand and supply of corn and show the intercept values, equilibrium quantity

and price.


c. What is the total revenue derived in this market?


1
Expert's answer
2021-09-12T19:31:47-0400

a)

"P = 100 – \frac{1}{2}Q\\
P = 10 + \frac{13}{10}Q.\\100 – \frac{1}{2}Q
= 10 + \frac{13}{10}Q.\\90=1.3Q+0.5Q\\90=1.8Q\\Q=50\\P=100-0.5(50)\\=100-25\\=75"


b)



c)

"Total\space revenue=price×quantity\\=75×50\\=3750"